Biological Tissues are typically made up of 98% water.Given that the density of water is 1.0x10^-3 kg/m3. Estimate the mass of a heart of a young man.

b: mass of a cell with a diameter of 0.5 x 10^-6 m


1
Expert's answer
2021-12-12T16:42:34-0500

The mass of a heart of a young man is 300 g.

The mass of the cell is


"m=\rho ·\frac 43\pi r^3=5.24·10^{-16}\text{ kg}."